# Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

### v6

@@ 3.2. Challenges in Responsible Generative AI
We group GenAI challenges into technical and regulatory domains.
3.2.1. Technical Challenges
@@
Data-Related Challenges
AI models rely on diverse and representative training data to achieve equitable performance across demographic groups and application contexts. Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities. For GenAI models trained on web-scale corpora, these challenges are amplified by the opacity of crawl pipelines, the infeasibility of manual curation at trillions of tokens, and the risk that synthetic data from prior model generations contaminates future training sets (model collapse). Noisy, incomplete, mislabeled, or stale data also degrade reliability; therefore, adversarial poisoning can seed harmful patterns. Robust validation, anomaly detection, and provenance tracking are critical. Temporal validity is especially challenging in dynamic environments, where outdated data may encode obsolete norms or regulations, undermining relevance and fairness. ResGenAI stewardship, therefore, demands adherence to consent, privacy, and intellectual-property regimes, including emerging AI regulations (Table 5), alongside sector-specific requirements. Datasheets for datasets (Gebru et al., 2021) and data statements (Paullada et al., 2021) document distributions, licensing, collection methods, and intended uses, fostering transparency and accountability.
@@
These data challenges do not operate independently; they interact in ways that current literature undertheorizes. Representativeness goals can conflict with privacy regimes: expanding demographic coverage may require collecting sensitive attributes that GDPR restricts. Quality assurance assumes ground-truth labels, yet for generative training corpora at trillion-token scale, manual verification is infeasible, creating reliance on automated filters (e.g., toxicity classifiers) that themselves embed biases. Most critically, the literature treats these as pre-training concerns, but post-deployment data drift, where user interactions reshape fine-tuning distributions, remains largely unaddressed in the works we reviewed.
Model- and System-Related Challenges
@@
Model- and system-level risks span explainability, bias amplification, robustness under adversarial conditions, calibration, and post-deployment drift. LLMs are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings. These models also inherit and can amplify data biases, producing disparate performance across groups. Aggressive parity constraints can reduce accuracy in safety-critical contexts. In generative settings, these biases manifest not only as allocational harms but as representational harms in generated text and images (e.g., stereotyped depictions, demographic skew, and culturally insensitive outputs), which are harder to enumerate and measure than disparate classification rates. Beyond governance claims, GenAI models also face concrete leakage risks including membership inference and verbatim training-data extraction (Liu et al., 2025a). Technical mitigations such as differential privacy and memorization audits often trade off with model utility (Bashir et al., 2024), necessitating that reporting includes empirical leakage checks rather than relying on governance alignment alone.
Current static, one-shot benchmarks rarely capture these dynamics. Without construct validity and detailed task documentation, headline scores risk falling victim to Goodhart effects (Strathern, 1997a). Safety claims should instead be framed as assurance cases with traceable evidence (e.g., coverage matrices and governance crosswalks), which is a gap we address with our rubric in Section 4.2.1.
3.2.2. Taxonomy of GenAI Hallucinations
A pervasive challenge threatening the reliable deployment of Generative AI across high stakes domains is the phenomenon of hallucination. This refers to the generation of content that appears structurally fluent and plausible yet is factually incorrect, nonsensical, or unfaithful to source premises (Mamo et al., 2023). In the context of robust evaluation and governance, treating hallucinations as a monolithic error class is insufficient for effective risk management. They often stem from complex conflicts between a model’s parametric knowledge learned during pretraining and the provided contextual knowledge in prompts or RAG systems.
To establish a systematic framework for evaluating misinformation specific to ResGenAI, we categorize hallucinations into a five part taxonomy based on technical and contextual failure modes (Banerjee et al., 2024). This taxonomy, visualized in Fig. 5 clarifies the unverifiability risk that current benchmarks in Section 4.2.1 and sectoral applications in Section 5 must address. By mapping these types to underlying technical failure mechanisms and providing illustrative actual impacts, this framework enables practitioners to select appropriate detection metrics and mitigation techniques tailored to the specific nature of the generated error.
@@ 3.2.2. Taxonomy of GenAI Hallucinations
Figure 5. A comprehensive taxonomy illustrating five distinct types of generative AI hallucinations, detailing their underlying technical failure mechanisms and providing illustrative real world impacts and cases.
3.2.3. Regulatory and Broader Sociotechnical Challenges
Regulatory challenges concern the design, implementation, and enforcement of frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions. User behavior and deployment context mediate how risks materialize.
@@ 3.2.3. Regulatory and Broader Sociotechnical Challenges
The EU AI Act proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024). Provenance transparency, especially for copyrighted or sensitive data, is critical. Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale, creating tension between compliance and performance (Slattery et al., 2024); dynamic risks such as election misinformation and deepfakes further require evolving standards. Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified. Over-reliance on assistive GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts. Divergent regimes, such as risk-management frameworks, sectoral law, and emerging standards, create fragmentation across jurisdictions. Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments. Inconsistent processes for dataset governance, evaluation, deployment, and monitoring further exacerbate bias and privacy risks (Slattery et al., 2024); pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
3.3. Emergent Threats
@@
While Section 3.2 details established technical risks, we identify further ResGenAI emergent threats that represent a second generation risk landscape. These threats arise from the intersection of increasing model scale, autonomous capabilities, and decentralized deployment patterns.
3.3.1. Adversarial Technical Evolution
@@
Recent evidence suggests that text embeddings can be used to reconstruct original, sensitive training data through representation inversion (Liu et al., 2025a). Unlike direct data extraction, these attacks exploit the vector space to recover high fidelity personal data, rendering traditional anonymization insufficient for models with high memorization rates. Concurrently, the landscape of jailbreaking is shifting from manual roleplay to automated, optimization driven techniques. Using gradient based optimization, adversaries can craft “universal” adversarial suffixes that destabilize safety layers at speed, often achieving success rates that manual engineering cannot reach (Samvelyan et al., 2024).
3.3.2. Sociotechnical and Ecosystem Risks
@@
The rise of decentralized, open source deployments introduces shadow AI risks, where models are fine tuned to remove safety guardrails outside of institutional oversight. Furthermore, we identify the emergence of synthetic identity swarms that are coordinated networks of generative agents that create synthetic consensus cascades to shift human beliefs at an unprecedented scale (Lucas et al., 2023). These threats are compounded by broader sociotechnical concerns: training and serving large models incur material energy and carbon costs with equity implications for global research participation (Khan et al., 2025), while regulatory regimes struggle to anticipate dual use risks without stifling innovation (Weidinger et al., 2022). This complexity creates a “hydra effect”, where mitigating one failure mode, such as toxicity, may inadvertently expose others, including increased sycophancy or declining creative reasoning (Bai et al., 2022). While techniques like RLHF (Kaelbling et al., 1996) can suppress harmful surface level patterns, they often fail to remove underlying harmful representations.
3.3.3. Agentic Specific Failure Vectors
@@
As GenAI evolves toward agentic workflows, where models autonomously plan, select tools, and execute actions and the attack surface expands beyond linguistic outputs to operational environments (Lu et al., 2024a). Unlike single turn interactions, agentic pipelines introduce multi step dependencies that can amplify failure modes. Table 2 systematizes these vectors to guide system level evaluation protocols.
@@
Table 2. Agentic Specific Failure Vectors: Risks in Autonomous Tool Use and Planning
@@
Tool Misuse Unintended or unauthorized API calls resulting from prompt injection or planning errors (Lu et al., 2024a). Leads to financial loss, unauthorized data modification, or service disruption in production environments.
@@
Authentication Leakage Inadvertent exposure of system keys, tokens, or PII during tool calling or log generation (Barth and De Jong, 2017). Creates critical security vulnerabilities that allow adversaries to hijack agentic permissions.
Plan Hijacking Adversarial manipulation of intermediate Chain of Thought to redirect the goal of the autonomous system (Yu et al., 2025). Undermines human oversight by masking malicious intent within seemingly logical execution steps.
This systemic fragility motivates the use of proportionate, auditable controls and transparent evidence. In Section 4.2, we operationalize this via our comparative rubric and provide a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
@@ 4.1. Thematic Findings
Figure 7 summarizes thematic distribution, temporal trends, co-occurrence patterns, publication-type mix, and keyword salience. Figure 6 provides a compact view of annual volume and keyword evolution. Appendix Table 2 list representative studies included in this review. Our thematic synthesis highlights that ResGenRAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF and others (Table 5); (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture. Our temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2026. These themes frequently co-occur, for example where explainability is invoked to mitigate fairness concerns, or where privacy and governance frameworks intersect in deployment contexts.
@@ 4.2. Benchmarks: Comparative Assessment
This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in Tables 4 and 6.
@@ 4.2.1. Comparative Assessment Protocol
Rubric design.
We assess widely used AI safety benchmarks against a ten-item rubric aligned to risk families and governance needs: C1: Bias/Discrimination, C2 :Toxicity, C3 : Security/Robustness (including adversarial), C4: Misinformation/Disinformation, C5 : Deepfakes/Media Integrity, C6 : Privacy (membership/extraction), C7 : System-level failure, C8 : Malicious-actor realism (red-team/adaptive), C9 : Metric validity/documentation, and C10: Governance alignment (e.g., NIST AI RMF, EU AI Act). The ten criteria (also given in Appendix A.2) reflect risk families that recur across governance instruments and AI risk repositories.
Scoring procedure and reliability.
Each benchmark was scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, $1$ = partial, $2$ = strong) by two independent reviewers, following the codebook in Appendix A.2. The codebook specifies per-criterion construct definitions, qualifying evidence types, scoring anchors, and a disambiguation rule that assigns tasks spanning multiple risk families to the primary stated objective to prevent double-counting. Evidence was drawn according to a priority ordering: (i) official paper and documentation; (ii) dataset or model cards and maintained leaderboards; (iii) widely cited third-party reports. Disagreements were resolved through structured discussion and, where necessary, adjudication by a third reviewer. Equal weights are used by default. The normalised score is
@@
This yields values in $[0,1]$, where 0 indicates no coverage and $1$ indicates full coverage across all criteria. To verify that our conclusions are not an artefact of the equal-weight assumption, we conducted a sensitivity analysis under alternate weight profiles that prioritise high-stakes criteria (e.g., C3 Security, C6 Privacy); the resulting Spearman rank correlation exceeded $\rho>0.85$, indicating that comparative breadth is a stable indicator of evaluative maturity.
Table 3. Mapping benchmarks to risk surface areas. $\checkmark$ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation). This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
Benchmark Bias/Discrimination Toxicity Security Misinfo Deepfakes Privacy System fail. Malicious
HarmBench (Mazeika et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SALAD-Bench (Li et al., 2024a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$
ALERT (Tedeschi et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Risk Taxonomy (Cui et al., 2024) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
TrustGPT (Huang et al., 2023b) $\checkmark$ $\checkmark$
RealToxicityPrompts (Gehman et al., 2020) $\checkmark$
DecodingTrust (Wang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
SafetyBench (Zhang et al., 2024) $\checkmark$ $\checkmark$ $\triangle$
MM-SafetyBench (Liu et al., 2024) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\checkmark$
XSTest (Röttger et al., 2024) $\triangle$ $\checkmark$ $\triangle$
Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\checkmark$ $\triangle$ $\checkmark$ $\checkmark$
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
HELM (Liang et al., 2023) $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$
SHIELD(Shi et al., 2024) $\triangle$ $\triangle$ $\checkmark$
BIG-bench (bench authors, 2023) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
PrivLM-Bench (Li et al., 2024b) $\checkmark$
PrivacyLens (Shao and et al., 2024) $\triangle$ $\checkmark$
HumaniBench (Raza et al., 2025) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Table 4. Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong). Scores use equal weights and are normalized to $[0,1]$. Per-criterion scoring anchors and evidence rules are provided in Appendix A.2. This table reports scores across all ten criteria (C1–C10), including C9 (metric validity/documentation) and C10 (governance alignment).
Benchmark C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 Score
@@
SafetyBench (Zhang et al., 2024) 1 1 1 0 0 0 0 0 1 1 0.25
MM–SafetyBench (Liu et al., 2024) 1 1 1 0 1 0 0 0 1 1 0.30
SALAD Bench (Li et al., 2024a) 1 1 1 0 0 0 0 0 1 1 0.25
HarmBench (Mazeika et al., 2024) 0 1 1 0 0 0 0 1 1 0 0.20
Rainbow Teaming (Samvelyan et al., 2024) 1 1 2 1 0 0 1 2 1 1 0.50
ALERT (Tedeschi et al., 2024) 0 1 1 0 0 0 0 2 1 1 0.30
HELM (Liang et al., 2023) 1 1 1 1 0 0 0 0 1 1 0.30
PrivLM Bench (Li et al., 2024b) 0 0 0 0 0 2 0 0 1 1 0.20
PrivacyLens (Shao and et al., 2024) 0 0 0 0 0 2 0 0 1 0 0.15
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) 1 1 1 1 0 0 0 1 1 1 0.35
SHIELD (Shi et al., 2024) 0 0 1 0 2 0 0 0 1 0 0.20
MFG for GenAI (Foundation, 2024) 1 0 1 1 0 0 0 1 0 0 0.20
XSTest (Röttger et al., 2024) 0 0 0 0 0 2 0 1 1 0 0.20
DecodingTrust (Wang et al., 2023a) 1 1 1 0 0 1 0 1 1 1 0.35
TrustLLM (Huang et al., 2024b) 1 1 0 1 0 1 1 0 1 0 0.30
BIG bench (bench authors, 2023) 1 1 1 0 0 0 1 0 0 0 0.25
Risk Taxonomy (Cui et al., 2024) 1 1 1 0 0 1 0 1 1 2 0.40
TrustGPT (Huang et al., 2023b) 1 1 0 0 0 0 0 0 1 0 0.20
HumaniBench (Raza et al., 2025) 2 1 0 1 0 1 1 0 1 1 0.40
Criteria (C1 to C10) C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment. Normalization (equal weights): $S(B)=\frac{\sum_{i=1}^{10}s_{i}(B)}{20}$ with $s_{i}\in\{0,1,2\}$.
@@
Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (ISO/IEC, 2021a): technical report on AI and data quality. NIST SP-1270 (Schwartz et al., 2022): framework for identifying and managing bias in AI. AI4ALL (AI4ALL, [n. d.]): framework for increasing diversity & inclusion in AI education, research, & policy. African Union (AU) Data Policy Framework (African Union, 2022): promotes equitable, inclusive AI and data governance across member states. Singapore IMDA & PDPC’s AI Governance Framework (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): addresses fairness, human-centricity, and non-discrimination in AI decision-making. Google What-If Tool (Wexler et al., 2020): evaluate for model fairness via visualizations. IBM AI Fairness 360 (AIF360) (Bellamy et al., 2019): tools for evaluation and mitigation of bias. Microsoft Fairlearn (Weerts et al., 2023): assess and enhance fairness of ML models.
@@
Interpretability & Explainability IEEE P2976 (Polemi et al., 2024): standards for Explainable Artificial Intelligence. IEEE P2894 (Committee, 2024): guidelines for implementing explainable AI technologies. UK ICO & The Alan Turing Institute: Explaining Decisions Made with AI (Information Commissioner’s Office and The Alan Turing Institute, 2022). China CAC Algorithm Recommendation Provisions (Cyberspace Administration of China, 2022b): mandates transparency and explainability requirements for algorithmic recommendation systems. Learning Interpretability Tool (LIT) (Google PAIR, [n. d.]): model interpretability and explainability. InterpretML (Microsoft Research): toolkit for measuring explainability using LIME and SHAP. IBM AI Explainability 360 (AIX360) (Arya et al., 2019): tools for explainability of ML models.
@@
Transparency & Accountability U.S. Government Accountability Office: AI Accountability Framework (U.S. Government Accountability Office, 2021). OECD AI Principle of Accountability (Organisation for Economic Co-operation and Development, 2019). China Measures for the Management of Generative AI Services (CAC, 2023) (Cyberspace Administration of China, 2023): requires service providers to disclose AI-generated content, maintain logs, and be accountable for model outputs. Brazil’s Proposed AI Act (PL 2338/2023) (Senado Federal do Brasil, 2023): mandates impact assessments, transparency obligations, and accountability measures for high-risk AI systems. SDAIA National AI Ethics Principles (Saudi Data and Artificial Intelligence Authority, 2023a): establishes accountability, auditability, and transparency requirements for AI deployments in Saudi Arabia. Evidently (Evidently AI, [n. d.]): evaluate, test and monitor ML models. MLBench (MLBench Team, [n. d.]): improve model transparency, reproducibility, and robustness. AI Verify (Infocomm Media Development Authority ([n. d.]), IMDA): Singapore’s IMDA-developed AI governance testing framework.

### v7

@@ 3.2. Challenges in Responsible Generative AI
Technical Challenges
@@
Data-Related Challenges. For GenAI models trained on web-scale corpora, data-related challenges are amplified by the opacity of crawl pipelines, the infeasibility of manual curation at trillions of tokens, and the risk that synthetic data from prior model generations contaminates future training sets (model collapse). Noisy, incomplete, mislabeled, or stale data degrade reliability; adversarial poisoning can seed harmful patterns; and temporal validity matters where outdated data may encode obsolete norms. ResGenAI stewardship demands adherence to consent, privacy, and intellectual-property regimes, including emerging AI regulations (Table 4), alongside sector-specific requirements. These challenges interact in ways that current literature undertheorizes. Representativeness goals can conflict with privacy regimes: expanding demographic coverage may require collecting sensitive attributes that GDPR restricts. Quality assurance assumes ground-truth labels, yet for trillion-token corpora manual verification is infeasible, creating reliance on automated filters that themselves embed biases. Most critically, the literature treats these as pre-training concerns, but post-deployment data drift—where user interactions reshape fine-tuning distributions—remains largely unaddressed.
@@
Model and System Related Challenges. LLMs are often treated as “black boxes” and can amplify data biases, producing disparate performance across groups. In generative settings, these biases manifest not only as allocational harms but as representational harms in generated text and images (e.g., stereotyped depictions, demographic skew), which are harder to enumerate and measure than disparate classification rates. GenAI models also face concrete leakage risks including membership inference and verbatim training-data extraction (Liu et al., 2025b). Technical mitigations such as differential privacy and memorization audits often trade off with model utility (Bashir et al., 2024). Without construct validity and detailed task documentation, headline scores risk Goodhart effects (Strathern, 1997). Safety claims should instead be framed as assurance cases with traceable evidence (e.g., coverage matrices and governance crosswalks)—a gap we address with our rubric in Section 4.2.1.
@@
Figure 4. Types of GenAI hallucinations, detailing their underlying technical failure mechanisms and illustrative real-world impacts.
GenAI Hallucinations
A pervasive challenge for reliable GenAI deployment is hallucination: generation of content that appears fluent and plausible but is factually incorrect, nonsensical, or unfaithful to source premises (Mamo et al., 2023). Treating hallucinations as a monolithic error class is insufficient for effective risk management. We categorize hallucinations into a five-part taxonomy based on technical and contextual failure modes (Banerjee et al., 2024), visualized in Fig. 4, to clarify the unverifiability risk that current benchmarks in Section 4.2.1 and sectoral applications in Section 5 must address.
Regulatory and Broader Sociotechnical Challenges
@@
Regulatory challenges concern designing, implementing, and enforcing frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions. Provenance transparency, especially for copyrighted or sensitive data, is critical, and legal expectations for interpretability in high-risk settings meet practical limits as models scale. Over-reliance on assistive GenAI can mask failure modes; documenting escalation paths to human review is essential in high-stakes contexts. Divergent regimes—risk-management frameworks, sectoral law, and emerging standards—create fragmentation across jurisdictions. Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments.
Emergent Threats
@@
While Section 3.2 details established technical risks, we identify further ResGenAI emergent threats representing a second-generation risk landscape, arising from the intersection of increasing model scale, autonomous capabilities, and decentralized deployment.
@@
Adversarial Technical Evolution. Recent evidence shows that text embeddings can reconstruct sensitive training data through representation inversion (Liu et al., 2025b), rendering traditional anonymization insufficient for models with high memorization rates. Concurrently, jailbreaking is shifting from manual role-play to automated, optimization-driven techniques: gradient-based optimization can craft “universal” adversarial suffixes that destabilize safety layers at speed and scale.
@@
Sociotechnical and Ecosystem Risks. Decentralized, open-source deployments introduce shadow AI risks, where models are fine-tuned to remove safety guardrails outside institutional oversight. We further identify synthetic identity swarms—coordinated networks of generative agents creating synthetic consensus cascades to shift human beliefs at unprecedented scale. These threats are compounded by material energy and carbon costs with equity implications for global research participation (Khan et al., 2025), and by regulatory struggles to anticipate dual-use risks without stifling innovation. This creates a “hydra effect”: mitigating one failure mode (e.g., toxicity) may expose others, including increased sycophancy or declining creative reasoning (Bai et al., 2022). While RLHF (Kaelbling et al., 1996) can suppress harmful surface-level patterns, it often fails to remove underlying harmful representations.
@@
Agentic-Specific Failure Vectors. As GenAI evolves toward agentic workflows—where models autonomously plan, select tools, and execute actions—the attack surface expands beyond linguistic outputs to operational environments (Lu et al., 2024a). Unlike single-turn interactions, agentic pipelines introduce multi-step dependencies that amplify failure modes. Table 2 systematizes these vectors; in Section 4.2, we operationalize this via our comparative rubric and governance crosswalk mapping risks to evaluative evidence.
@@
Table 2. Agentic-Specific Failure Vectors: Risks in Autonomous Tool Use and Planning
@@
Tool Misuse Unintended or unauthorized API calls from prompt injection or planning errors (Lu et al., 2024a). Financial loss, unauthorized data modification, or service disruption.
@@
Authentication Leakage Inadvertent exposure of keys, tokens, or PII during tool calling or log generation (Barth and De Jong, 2017). Critical security vulnerabilities allowing adversaries to hijack agentic permissions.
Plan Hijacking Adversarial manipulation of intermediate Chain of Thought to redirect autonomous system goals (Yu et al., 2025). Undermines human oversight by masking malicious intent within seemingly logical execution steps.
@@ 4.1. Thematic Findings
Figure 6 summarizes thematic distribution, temporal trends, co-occurrence patterns, publication-type mix, and keyword salience. Figure 5 provides a compact view of annual volume and keyword evolution. Appendix Table 2 lists representative studies included in this review. Our thematic synthesis highlights that ResGenRAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF and others (Table 4); (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture. Our temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2026.
@@ 4.2. Benchmarks: Comparative Assessment
This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in Tables 3 and 5.
@@ 4.2.1. Comparative Assessment Protocol
Rubric design: We assess widely used AI safety benchmarks against a ten-item rubric aligned to risk families and governance needs: C1: Bias/Discrimination, C2 : Toxicity, C3 : Security/Robustness (including adversarial), C4: Misinformation/Disinformation, C5 : Deepfakes/Media Integrity, C6 : Privacy (membership/extraction), C7 : System-level failure, C8 : Malicious-actor realism (red-team/adaptive), C9 : Metric validity/documentation, and C10: Governance alignment (e.g., NIST AI RMF, EU AI Act). The ten criteria (also given in Appendix 1.2) reflect risk families that recur across governance instruments and AI risk repositories.
Scoring procedure and reliability: Each benchmark was scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, $1$ = partial, $2$ = strong) by two independent reviewers, following the codebook in Appendix 1.2. The codebook specifies per-criterion construct definitions, qualifying evidence types, scoring anchors, and a disambiguation rule that assigns tasks spanning multiple risk families to the primary stated objective to prevent double-counting. Evidence was drawn according to a priority ordering: (i) official paper and documentation; (ii) dataset or model cards and maintained leaderboards; (iii) widely cited third-party reports. Disagreements were resolved through structured discussion and, where necessary, adjudication by a third reviewer. Equal weights are used by default. The normalized score is
@@
This yields values in $[0,1]$, where 0 indicates no coverage and $1$ indicates full coverage across all criteria. To verify that our conclusions are not an artifact of the equal-weight assumption, we conducted a sensitivity analysis under alternate weight profiles that prioritize high-stakes criteria (e.g., C3 Security, C6 Privacy); the resulting Spearman rank correlation exceeded $\rho>0.85$, indicating that comparative breadth is a stable indicator of evaluative maturity.
Table 3. Comparative assessment rubric for AI safety benchmarks. Criteria C1–C8 double as a risk-surface map: 0 = no coverage, 1 = partial, 2 = strong. Scores use equal weights and are normalized to $[0,1]$. Per-criterion scoring anchors and evidence rules are provided in Appendix 1.2.
Risk Surface (C1–C8) Quality
Benchmark
C1 Bias/Discrim.
C2 Toxicity
C3 Security/Rob.
C4 Misinfo.
C5 Deepfakes
C6 Privacy
C7 Sys. Failure
C8 Malicious
C9 Metric Val.
C10 Gov. Align.
Score
Rainbow Teaming (Samvelyan et al., 2024) 1 1 2 1 0 0 1 2 1 1 0.50
Risk Taxonomy (Cui et al., 2024) 1 1 1 1 0 1 0 1 1 2 0.45
HumaniBench (Raza et al., 2025) 2 1 0 1 0 1 1 0 1 1 0.40
DecodingTrust (Wang et al., 2023) 1 1 1 1 0 1 0 1 1 1 0.40
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) 1 1 1 1 0 0 0 1 1 1 0.35
SALAD-Bench (Li et al., 2024b) 1 1 1 1 0 0 0 1 1 1 0.35
ALERT (Tedeschi et al., 2024) 1 1 1 0 0 0 0 2 1 1 0.35
MM-SafetyBench (Liu et al., 2024) 1 1 1 1 1 0 0 0 1 1 0.35
HELM (Liang et al., 2023) 1 1 1 1 0 0 0 0 1 1 0.30
TrustLLM (Huang et al., 2024b) 1 1 0 1 0 1 1 0 1 0 0.30
HarmBench (Mazeika et al., 2024) 0 1 1 1 0 0 0 1 1 0 0.25
SafetyBench (Zhang et al., 2024) 1 1 1 0 0 0 0 0 1 1 0.25
XSTest (Röttger et al., 2024) 0 1 0 0 0 2 0 1 1 0 0.25
SHIELD (Shi et al., 2024) 1 0 1 0 2 0 0 0 1 0 0.25
BIG-bench (authors, 2023) 1 1 1 1 0 0 1 0 0 0 0.25
PrivLM-Bench (Li et al., 2024a) 0 0 0 0 0 2 0 0 1 1 0.20
MFG for GenAI (Foundation, 2024) 1 0 1 1 0 0 0 1 0 0 0.20
PrivacyLens (Shao and et al., 2024) 1 0 0 0 0 2 0 0 1 0 0.20
@@
TrustGPT (Huang et al., 2023b) 1 1 0 0 0 0 0 0 1 0 0.15
Scoring: $s_{i}\!\in\!\{0,1,2\}$; $S(B)=\sum s_{i}/20$. C1–C8 also serve as the risk-surface crosswalk (0 $\hat{=}$ no coverage, 1 $\hat{=}$ partial, 2 $\hat{=}$ full). C9: Metric validity & documentation; C10: Governance alignment.
@@
Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (ISO/IEC, 2021a): technical report on AI and data quality. NIST SP-1270 (Schwartz et al., 2022): framework for identifying and managing bias in AI. AI4ALL (AI4ALL, ): framework for increasing diversity & inclusion in AI education, research, & policy. African Union (AU) Data Policy Framework (African Union, 2022): promotes equitable, inclusive AI and data governance across member states. Singapore IMDA & PDPC’s AI Governance Framework (Infocomm Media Development Authority (IMDA) and Personal Data Protection Commission (PDPC), 2020): addresses fairness, human-centricity, and non-discrimination in AI decision-making. Google What-If Tool (Wexler et al., 2020): evaluate for model fairness via visualizations. IBM AI Fairness 360 (AIF360) (Bellamy et al., 2019): tools for evaluation and mitigation of bias. Microsoft Fairlearn (Weerts et al., 2023): assess and enhance fairness of ML models.
@@
Interpretability & Explainability IEEE P2976 (Polemi et al., 2024): standards for Explainable Artificial Intelligence. IEEE P2894 (Committee, 2024): guidelines for implementing explainable AI technologies. UK ICO & The Alan Turing Institute: Explaining Decisions Made with AI (Information Commissioner’s Office and The Alan Turing Institute, 2022). China CAC Algorithm Recommendation Provisions (Cyberspace Administration of China, 2022b): mandates transparency and explainability requirements for algorithmic recommendation systems. Learning Interpretability Tool (LIT) (Google PAIR, ): model interpretability. InterpretML: toolkit for explainability IBM AI Explainability 360 (AIX360) (Arya et al., 2019): tools for explainability.
@@
Transparency & Accountability U.S. Government Accountability Office: AI Accountability Framework (U.S. Government Accountability Office, 2021). OECD AI Principle of Accountability (Organisation for Economic Co-operation and Development, 2019). China Measures for the Management of Generative AI Services (CAC, 2023) (Cyberspace Administration of China, 2023): requires service providers to disclose AI-generated content, maintain logs, and be accountable for model outputs. Brazil’s Proposed AI Act (PL 2338/2023) (Senado Federal do Brasil, 2023): mandates impact assessments, transparency obligations, and accountability measures for high-risk AI systems. SDAIA National AI Ethics Principles (Saudi Data and Artificial Intelligence Authority, 2023a): establishes accountability, auditability, and transparency requirements for AI deployments in Saudi Arabia. Evidently (Evidently AI, ): evaluate, test and monitor ML models. MLBench (MLBench Team, ): improve model transparency, reproducibility, and robustness. AI Verify (Infocomm Media Development Authority (IMDA), ): Singapore’s IMDA-developed AI governance testing framework.